OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.A(5,5), M=3

OpenStudy (anonymous):

Start with point-slope form: \[y-5=3(x-5)\] distribute: \[y-5=3x-15\] subtract 3x: \[y-3x-5=-15\] add 5 : \[y-3x=-10 \rightarrow -3x+y=-10\] and multiply by -1: \[3x-y=10\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points.E(-2, 2), F(5, 1)

OpenStudy (anonymous):

Same thing, except the slope is determined by: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

Then: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

thanks that helped alot